# 4D Millimeter-Wave Radar in Autonomous Driving: A Survey

### v3

@@ IV Learning-based Radar Data Enhancement
Figure 8: Two dominant pipelines in learning-based radar data enhancement.
In the contemporary landscape of mmwave radar technology, two dominant pipelines have emerged based on learning methods. We categorize them as "Reconstructor" and "Detector". Their main difference are shown in Fig. 8.
IV-A Reconstructor
@@
Reconstructor focuses on improving the resolution of previously acquired radar point clouds. This approach is dedicated to the post-processing enhancement of data fidelity, thereby increasing the detail and usefulness of the radar imagery.
@@
The related work is primarily inspired by the reconstruction of LiDAR point clouds, such as the PointNet [37] structure in [38], which in turn inspired the works of Sun et al. [39] [36] [40]. Although these studies require data observed from different viewpoints of the target, which may constrain their immediate integration into autonomous driving systems, the underlying principles of their model architectures offer valuable insights. Specifically, the employment of a conditional Generative Adversarial Network (GAN) architecture facilitates the concurrent training of generator and discriminator networks, as detailed in [39]. Moreover, the innovative two-stage point cloud generation process, which incorporates a loss function that synergistically combines Chamfer Distance (CD) and Earth Mover’s Distance (EMD) metrics, is elucidated in [36]. When confronted with coarse and sparse input point clouds, the methodologies proposed by Sun et al. have been validated to substantially surpass existing techniques, including PointNet[37], PointNet++[41] and PCN[38]. The robustness of these methods underscores their potential to significantly enhance the accuracy and reliability of point cloud reconstruction, even in the presence of suboptimal input radar data.
IV-B Detector
Meanwhile, Detector signifies a substantial departure from traditional processing techniques. This innovative approach leverages neural networks to engage directly with raw radar data such as RD maps or 4D tensors, which circumvents conventional techniques such as CFAR or DBF, potentially leading to more efficient and robust detection capabilities in real-time applications.
Brodeski et al. [16] pioneer the application of Convolutional Neural Networks (CNNs) to RD maps for the detection and localization of multiple objects called DRD(deep radar detection) net. They reconceptualize the challenge of target detection within RD maps as a segmentation task and adopt a model structure analogous to the 2D-U-Net[42], renowned for its efficacy in medical image segmentation. Confronted with the scarcity of well-annotated RD map datasets, they devise a strategy to extract labeled radar data from the calibration process conducted within an anechoic chamber. Furthermore, they introduce a novel data augmentation method to enhance the diversity and volume of the dataset. Experiments of the DRD network demonstrate its capability to function in real-time, with inference times recorded at approximately 20ms. Notably, the DRD network has been shown to surpass classic methods in terms of detection accuracy and robustness. Though this work does not include real-world radar data with all the impairments that come along, the findings from this study unequivocally illustrate the considerable promise of neural network applications to radar complex data.
@@ IV-B Detector
IV-C Challenge
As noted in [27, 45], the diversity of mmWave radar hardware utilized across various datasets leads to significant variations in pre-CFAR data, making it challenging to construct a large-scale public dataset and benchmark. Furthermore, pre-CFAR data is far less intuitive than point cloud data, rendering the manual labeling process exceedingly laborious and difficult to achieve accuracy, hence producing high-quality supervised data for neural networks is problematic. Another avenue is the adoption of learning-based approaches for generating synthetic automotive radar scenes and data[46, 47]. However, as such methods are based on simulation, it depends on a world model that may contain inaccuracies. Additionally, the management of pre-CFAR data is associated with considerable computational and memory demands, necessitating a judicious balance between the requirements for computational power and the minimization of information loss.
@@ IV-C Challenge
V Perception Applications
Currently, the point cloud density of 4D mmWave radars has already attained a level comparable to that of low-beam LiDAR, with the added advantages of exhibiting superior robustness under low visibility and adverse weather conditions. Therefore, researchers have been attempting to transfer LiDAR point cloud processing models to 4D mmWave radars in various tasks, including target detection[53, 51, 54], trajectory tracking[55, 56], and scene flow prediction[57, 58], among others. Furthermore, as described in Section IV, pre-CFAR radar data encompasses a wealth of information, promoting some researchers to engage directly with RD maps or 4D tensors, bypassing point cloud generation tasks. Existing 4D mmWave radar point cloud and pre-CFAR data feature extraction methods in autonomous driving perception are summarized in Fig. 11, which will be detailed in this section.
In Section V-A, we review and analyze perception models for radar point clouds (RPC), which are primarily enhancements of LiDAR-based methodologies. The 4D mmWave radars branch of some fusion methods are also included. In Section V-B, we investigate the methods utilizing pre-CFAR radar data, including the range-frequency map, range-azimuth map, range-azimuth-Doppler cube, and 4D tensors. In Section V-C, the integration of 4D mmWave radars into multi-sensor fusion systems, as well as the effectiveness of such methods are present. In Section V-D, we discuss current challenges of this field.
TABLE II: SUMMARY OF 4D-RADAR PERCEPTION METHODS
@@ B. WITH PRE-CFAR DATA
V-A Point Cloud Feature Extraction
@@
Given the analogous nature of their data formats, it is clear that a significant number of RPC methodologies have their roots in LiDAR-based techniques. However, this transposition necessitates meticulous consideration of the inherent constraints associated with radar systems. These limitations encompass low-resolution representations, data sparsity, and inherent uncertainty within the data. Conversely, radar systems outperform in areas such as superior range resolution, velocity measurement capabilities, and the advantage of early target detection. Consequently, these unique characteristics necessitate the development of specifically tailored network designs.
Upon a thorough review of the literature, it appears that Meyer et al. [66] pioneer the adaptation of a CNN architecture, originally designed for camera-LiDAR fusion[75], for the processing of 4D mmWave radar point clouds and RGB images. To reconcile the discrepancy in data formats, they elect to omit the Doppler information inherent to 4D mmWave radar data, instead solely utilize the spatial coordinates and intensity values. Subsequently, they employ the resulting point clouds to generate Bird’s Eye View (BEV) images with six height maps and one density map to formulate three-dimensional (3D) object proposals. Remarkably, their fusion network demonstrates enhanced precision employing RPC instead of LiDAR point clouds, achieving an average precision (AP) of 61% on the Astyx dataset [25]. The authors postulate that this improvement may be attributable to the relatively sparse 16-beam configuration of the LiDAR sensor within the dataset. Although the authors recognise that this hypothesis need further exploration, the preliminary findings remain both intriguing and promising.
@@
Regarding subsequent research, Table II offers a comprehensive and succinct exposition of the field. Building upon the foundation of existing work, these investigations have masterfully exploited the distinctive attributes of 4D mmWave radar point clouds, encompassing elements like elevation, Doppler data, and Radar Cross Section (RCS) intensity. Moreover, they have ingeniously formulated strategies to address the inherent sparsity and irregular distribution of these data points, thereby advancing the field significantly.
@@
In light of the augmented elevation data presented by 4D mmWave radars, a distinct advancement over 3D mmWave radars, or the Doppler information, a feature not mirrored in LiDAR systems, most studies[66, 52, 50, 30, 67] have opted to reference the implicit structures like SECOND[76] and PointPillars[59]. These studies encode extra radar attributes directly, comparable to the conventional spatial coordinates $x,y,z$ in point clouds. Palffy et al.[30] demonstrate that the addition of elevation data, Doppler information, and RCS information respectively increase the 3D mean Average Precision (mAP) with 6.1%, 8.9% and 1.4%. However, the result of the proposed method (47.0% mAP) is still far inferior to the LiDAR detector on 64-beam LiDAR (62.1% mAP), indicating that there is still room for improvement in the optimization of 4D mmWave radar-based detection methods.Nevertheless, Zheng et al.[48] introduce a subtle yet impactful modification, by proposing the Radar PillarNet backbone, colloquially termed RPNet. This structure employs three separate linear layers, each with unshared weights, to extract spatial position, velocity, and intensity features, respectively. Subsequently, a BEV pseudo image is generated. Ablation studies have demonstrated that RPNet enhances the 3D mAP by 4.26%.
@@
Addressing the utilization of Doppler information, Tan et al. [50] delineate a widely-adopted yet efficacious technique. Recognizing that mmWave radars intrinsically measure the radial relative velocity of detected objects, and the motion of the vehicle itself results in different coordinate systems for multi-frame point clouds, they first calculate the vehicle’s velocity, followed by compensating and obtaining each point’s true velocity relative to the ground, which is often referred to as ’absolute velocity’. Moreover, the integration of Doppler information facilitates a more convenient and accurate accumulation of historical frame point clouds, thereby enhancing point cloud density. Yan et al. [49] further propose the Radar Feature Assisted Backbone. In this design, each point’s absolute and relative velocities, along with its reflectivity, are integrated into position embeddings. These embeddings are then multiplied with the self-attention reweighting map of point-wise features, thereby enhancing the exchange of information at the feature vector level in a trainable fashion. On the other hand, Pan et al. [55] introduce a ’detection by tracking’ strategy. This approach leverages velocity characteristics to achieve point-level motion segmentation and scene flow estimation. Subsequently, employing the classical DBSCAN clustering method suffices to surpass the tracking accuracy of established techniques like centerpoint[60] and AB3DMOT[77].
@@
In the domain of model backbone architecture, the incorporation of spatial attention mechanisms has been acknowledged as an effective strategy to address the sparsity and irregular distribution of point clouds. Xu et al.[52] implemented a self-attention mechanism to extract global information from the pillarized radar point cloud. Further advancing this concept, Shi et al.[63] augment different view features using a combination of self-attention and cross-attention mechanisms. While self-attention focuses on understanding the relationships within a single view, cross-attention extends this understanding across different views, thus enabling a more comprehensive and integrated feature representation. Yan et al.[49] take a different approach by utilizing the attention matrix inherent in the self-attention mechanism to differentiate and reweight foreground and background points and their respective features. They also introduce a binary classification auxiliary loss to aid the learning process. Additionally, several studies have used spatial attention to fuse multimodal sensor data, not only addressing the noise and sparsity in radar point clouds, but also leveraging the strengths of different sensor modalities. These methods will be further elaborated in Sect.V-C.
V-B Pre-CFAR Feature Extraction
@@
In the domain of millimeter-wave (mmWave) radar signal processing, techniques such as side lobe suppression and CFAR algorithms are integral for mitigating noise and minimizing false alarms. These methodologies facilitate the extraction of signal peaks, thereby reducing the volume of data and decreasing computational demands. However, a consequential drawback of this approach is the sparsity of radar point clouds, characterized by diminished resolution. Given the profound advancements in the realm of deep learning, particularly in the processing of dense image data, a pivot in research focuses towards Pre-CFAR data is observed, aiming to utilize more underlying hidden information.
@@
To our best knowledge, there are currently several datasets containing 4D mmWave radar Pre-CFAR data [31, 28, 27, 26, 79]. Notably, a subset of these datasets [27, 26, 79] is characterized by a relatively lower elevation resolution, exceeding 15 degrees. Consequently, our survey will predominantly focus on the methodologies employed within the high resolution datasets provided by [31, 28]. This section intends to elucidate the comprehensive pipeline and the enhancements tailored for optimizing the characteristics of 4D mmWave radar data.
@@
As discussed in Section II-A, the 4D radar signal processing workflow employs FFT methodologies on raw ADC data to discretely process the four dimensions: range, Doppler, azimuth, and elevation. This processing culminates in the generation of diverse data representations, including the Range-Doppler (RD) map, Range-Azimuth (RA) map, Range-Azimuth-Doppler (RAD) cube, and ultimately, a 4D tensor. From 2D maps to 4D tensors, the complexity of extracted features varies. The more dimensions of information contained in the data, the higher the memory and computation costs will be for the feature extraction process. The extracted features are generally aligned to the RA axis in BEV under polar coordinates or the XY axis in Cartesian coordinates, connecting to detection or segmentation heads, serving as the foundational elements for subsequent detection or segmentation operations.
Figure 13: Overview of RTNH[74] with 4D radar tensors
@@
For the 4D radar tensor, Paek et al. [74, 31] opt to further extract a sparse tensor aligned to the Cartesian coordinate system, subsequently utilizing 3D sparse convolution to extract multi-scale spatial features. Experiments have demonstrated that retaining only the top-5% elements with the highest power measurements can maintain detection accuracy while significantly enhancing processing speed. Furthermore, the elevation information included in 4D tensors is essential facing 3D but not BEV 2D object detection.
@@
Conversely, RA map data, inherently aligned with the polar coordinate system, is amenable to direct processing through dense feature extraction networks. The generated BEV features are then either converted into the Cartesian coordinate system via bi-linear interpolation [27] or utilized within polar-based detection frameworks [73].
@@
V-C Multi-Sensor Fusion Methods
@@
Exemplifying this approach, Meyer et al.[66] employ a 3D region proposal network to camera images and BEV images of the radar pointcloud, then meticulously crop, resize, and integrate multi-modal features for each proposed anchor. A subsequent study is performed by Cui et al.[53] with a novel self-supervised model adaptation block[80], which dynamically adapts the fusion of different modalities in accordance with the object properties. Besides, a FV map is generated from the 4D mmWave radar point clouds together with the BEV image. The presented method outperforms the former study[66] by up to 9.5% in 3D AP. The FV map effectively leverages the elevation information provided by 4D mmWave radars and achieves easier fusion with the monocular camera feature, balancing detection accuracy and computational efficiency.
Additionally, recent works such as RCFusion[48] and LXL[62] have advanced the integration of attention mechanisms for the fusion of image and 4D radar features. They begin by separately extracting BEV features from the camera and radar branches, then employ convolutional networks to create scale-consistent attention maps, effectively delineating the occupancy grid for target objects. The distinction lies in the fact that RCFusion[48] generates 2D attention maps in both the camera and radar branches, which are then multiplied with the BEV feature from the other modality. Conversely, LXL[62] solely utilizes the radar BEV feature to infer the 3D occupancy grid, which is then multiplied with the 3D image voxel features to achieve attention sampling of the image features.
@@
V-D Challenge
Current methodologies for 4D radar point cloud perception predominantly adapt established techniques from LiDAR applications, whereas approaches for pre-CFAR data often draw from the vision domain. Although the data formats exhibit similarities, the unique characteristics of mmWave radar data, specifically Doppler velocity and intensity information, warrant more focused attention for effective feature extraction. Moreover, pre-CFAR radar data characteristically contains a significantly higher ratio of background to actual objects (foreground)[71], a factor that complicates data interpretation and model training. The inherent noise within radar data further presents a substantial challenge for learning algorithms. The resilience of 4d radar models under out-of-distribution (OoD) conditions also remains inadequately explored and understood[31], which underscores the necessity for refined methodologies that can more accurately account for the distinct properties of mmWave radar data, thereby enhancing model robustness and performance in real-world scenarios.
VI Localization and Mapping Applications
In challenging environments where satellite-based positioning is unreliable or high-definition maps are absent, localization and mapping by perception sensors becomes indispensable. A body of relevant research has been conducted utilizing the emerging 4D mmWave radars. Given that radar point clouds are significantly less data-intensive than their tensor counterparts, and methodologies developed for LiDARs can be adapted to 4D mmWave radar point clouds with minimal modifications, the majority of localization and mapping studies preferentially employ radar point clouds as their input data instead of radar tensors. As Fig16 demonstrates, the unique Doppler velocity information contained within radar point clouds presents a notable opportunity for exploitation, which will be discussed in Section VI-A. Subsequently, both traditional and learning-based localization and mapping approaches will be introduced in Section VI-B and Section VI-C, respectively. And Section VI-D briefly introduces current challenges in 4D mmWave radar-based localization and mapping.
Figure 16: The taxonomy of 4D mmWave radar-based localization and mapping. Example images are from [81, 82, 29, 83, 84, 85, 86, 87, 88]
VI-A Doppler Information Utilizing
The Doppler information constitutes a significant advantage of 4D mmWave radar, especially in the context of localization and mapping applications. Nevertheless, the optimal exploitation of this information within the autonomous driving sphere remains a relatively open question. Generally speaking, the utilization of Doppler information can be categorized into several categories:
VI-A1 Ego-Velocity Estimation

### v4

@@ III Learning-based Radar Data Generation
Figure 7: Two dominant pipelines in learning-based radar data generation.
III-A Reconstructor
@@
Reconstructor methods focus on improving the resolution and detail of previously acquired radar point clouds. This approach is dedicated to the post-processing enhancement of data fidelity, thereby increasing the usefulness of the radar imagery.
@@
Much of the inspiration for these methods comes from the reconstruction of LiDAR point clouds. Notably, the PointNet structure [34] used in [35] has influenced subsequent studies by Sun et al. [36] [33] [37]. Although these studies require data from multiple viewpoints, which may constrain their immediate integration into autonomous driving systems, the underlying principles of their model architectures offer valuable insights. For example, they use a conditional Generative Adversarial Network (GAN) to train generator and discriminator networks concurrently, as detailed in [36]. Moreover, the innovative two-stage point cloud generation process, which incorporates a loss function that synergistically combines Chamfer Distance (CD) and Earth Mover’s Distance (EMD) metrics, is described in [33]. Sun et al.’s methods have shown significant improvements over existing techniques such as PointNet [34], PointNet++ [38], and PCN [35], particularly with coarse and sparse input point clouds. The robustness of these methods underscores their potential to enhance the accuracy and reliability of point cloud reconstruction, even with suboptimal radar data.
III-B Detector
Meanwhile, Detector approaches leverage neural networks to engage directly with raw radar data such as RD maps or 4D tensors, which circumvents conventional techniques such as CFAR or DBF, potentially leading to more efficient and robust detection capabilities in real-time applications.
Brodeski et al. [3] pioneer such frameworks and apply CNN-based image segmentation networks to RD maps for the detection and localization of multiple objects. Confronted with the scarcity of well-annotated RD map datasets, they devise a strategy to extract labeled radar data from the calibration process conducted within an anechoic chamber. Experiments of the DRD network demonstrate its capability to function in real-time, with inference times recorded at approximately 20ms. Notably, the DRD network has been shown to surpass classic methods in terms of detection accuracy and robustness. Though this work does not include real-world radar data with all the impairments that come along, the findings from this study unequivocally illustrate the considerable promise of neural network applications to radar complex data.
@@ III-B Detector
III-C Challenge
The development of Learning-based Radar Data Generation methods, particularly for 4D mmWave radar, is hindered by the scarcity of large, public datasets and benchmarks. As highlighted in recent studies [41, 42], the diversity of mmWave radar hardwares models complicates the standardization of pre-CFAR data andchallenges the creation of a comprehensive public dataset and benchmark. Furthermore, pre-CFAR data is far less intuitive than point cloud data, rendering manual annotation both laborious and prone to errors, thus complicating the production of high-quality supervised data for neural networks.
Another avenue is the adoption of learning-based approaches for generating synthetic automotive radar scenes and data[43, 44]. However, as such methods are based on simulation, it is limited by potential inaccuracies in the sensor and world model.
Moreover, the management of pre-CFAR data is associated with considerable computational and memory demands. While current methods employ lower-resolution radars to achieve real-time performance, accommodating higher resolution radars—which are essential for large outdoor scenes—requires further optimization of algorithmic efficiency.
@@ III-C Challenge
IV Perception Applications
Currently, the point cloud density of 4D mmWave radars has already attained a level comparable to that of low-beam LiDAR, with the added advantages of exhibiting superior robustness under low visibility and adverse weather conditions. Therefore, researchers have been attempting to transfer LiDAR point cloud processing models to 4D mmWave radars in various tasks, including target detection[48, 6, 49], trajectory tracking[50, 51], and scene flow prediction[52, 53], among others. Furthermore, as described in Section III, pre-CFAR radar data encompasses a wealth of information, promoting some researchers to engage directly with RD maps or 4D tensors, bypassing point cloud generation tasks. Existing 4D mmWave radar point cloud and pre-CFAR data feature extraction methods in autonomous driving perception are summarized in Fig. 10, which will be detailed in this section.
In Section IV-A, we review and analyze perception models for radar point clouds (RPC), which are primarily enhancements of LiDAR-based methodologies. The 4D mmWave radar branch of some fusion methods are also included. Section IV-B investigates the methods utilizing pre-CFAR radar data, including the range-frequency map, range-azimuth map, range-azimuth-Doppler cube, and 4D tensors. The integration of 4D mmWave radars into multi-modal fusion systems, as well as the effectiveness of such methods are present in Section IV-C. Finally in Section IV-D, we discuss current challenges of this field.
TABLE II: SUMMARY OF 4D MMWAVE RADAR PERCEPTION METHODS
@@ B. WITH PRE-CFAR DATA
IV-A Point Cloud Feature Extraction
@@
Given the analogous nature of their data formats, it is clear that a significant number of RPC methodologies originate from LiDAR-based techniques. Despite this similarity, this transposition requires careful consideration of the inherent constraints associated with radar systems. These limitations include low-resolution representations, data sparsity, and inherent uncertainty within the data. Conversely, radar systems outperform in areas such as superior range resolution, velocity measurement capabilities, and early target detection. Consequently, these unique characteristics necessitate the development of specifically tailored network designs.
@@
A comprehensive and succinct overview of recent advances in this field is presented in Table II. These investigations have masterfully exploited the distinctive attributes of 4D mmWave radar point clouds, encompassing elements like elevation, Doppler data, and Radar Cross Section (RCS) intensity. Moreover, they have ingeniously formulated strategies to address the inherent sparsity and irregular distribution of these data points, thereby advancing the field significantly.
IV-A1 Distinctive Information
4D mmWave radars provide a full three-dimensional view by measuring range, azimuth, and elevation of targets. Additionally, mmWave radars can measure the velocity of objects directly through the Doppler effect, a feature that distinguishes them from LiDAR systems. This combination of enhanced spatial data and velocity information makes 4D mmWave radars particularly valuable for Autonomous Driving tasks and presents unique opportunities for further studies.
@@ IV-A1 Distinctive Information
Most studies[62, 47, 46, 54, 63] have opted to reference the implicit structures like SECOND[73] and PointPillars[55]. These studies encode extra radar attributes directly, comparable to the conventional spatial coordinates $x,y,z$ in point clouds. Palffy et al.[54] demonstrate that the addition of elevation data, Doppler information, and RCS information respectively increase the 3D mean Average Precision (mAP) with 6.1%, 8.9% and 1.4%. However, the result of the proposed method (47.0% mAP) is still far inferior to the LiDAR detector on 64-beam LiDAR (62.1% mAP), indicating that there is still room for improvement in the optimization of 4D mmWave radar-based detection methods.Nevertheless, Zheng et al.[45] introduce a subtle yet impactful modification, by proposing the Radar PillarNet backbone, colloquially termed RPNet. This structure employs three separate linear layers, each with unshared weights, to extract spatial position, velocity, and intensity features, respectively. Subsequently, a BEV pseudo image is generated. Ablation studies have demonstrated that RPNet enhances the 3D mAP by 4.26%.
@@
Addressing the explicit utilization of Doppler information, Tan et al. [46] delineate a widely-adopted yet efficacious technique. Recognizing that mmWave radars intrinsically measure the radial relative velocity of detected objects, and the motion of the vehicle itself results in different coordinate systems for multi-frame point clouds, they first calculate the vehicle’s velocity, followed by compensating and obtaining each point’s true velocity relative to the ground, which is often referred to as ’absolute velocity’. Moreover, the integration of Doppler information facilitates a more convenient and accurate accumulation of historical frame point clouds, thereby enhancing point cloud density. Yan et al. [5] further propose the Radar Feature Assisted Backbone. In this design, each point’s absolute and relative velocities, along with its reflectivity, are integrated into position embeddings. These embeddings are then multiplied with the self-attention reweighting map of point-wise features, thereby enhancing the exchange of information at the feature vector level in a trainable fashion. On the other hand, Pan et al. [50] introduce a ’detection by tracking’ strategy. This approach leverages velocity characteristics to achieve point-level motion segmentation and scene flow estimation. Subsequently, employing the classical DBSCAN clustering method suffices to surpass the tracking accuracy of established techniques like centerpoint[56] and AB3DMOT[74].
IV-A2 sparsity and Irregularity
@@
In the domain of model backbone architecture, the incorporation of spatial attention mechanisms has been acknowledged as an effective strategy to address the sparsity and irregular distribution of point clouds. Xu et al.[47] implemented a self-attention mechanism to extract global information from the pillarized radar point cloud. Further advancing this concept, Shi et al.[59] augment different view features using a combination of self-attention and cross-attention mechanisms. While self-attention focuses on understanding the relationships within a single view, cross-attention extends this understanding across different views, thus enabling a more comprehensive and integrated feature representation. Yan et al.[5] take a different approach by utilizing the attention matrix inherent in the self-attention mechanism to differentiate and reweight foreground and background points and their respective features. They also introduce a binary classification auxiliary loss to aid the learning process. Additionally, several studies have used spatial attention to fuse multimodal sensor data, not only addressing the noise and sparsity in radar point clouds, but also leveraging the strengths of different sensor modalities. These methods will be further elaborated in Section IV-C.
IV-B Pre-CFAR Feature Extraction
@@
In millimeter-wave (mmWave) radar signal processing, side lobe suppression and CFAR algorithms play a crucial role in reducing noise and minimizing false alarms. These techniques help extract signal peaks, thereby reducing data volume and computational cost. However, a consequential drawback of this approach is the sparsity of radar point clouds, characterized by diminished resolution. Given the profound advancements in deep learning, particularly in the processing of dense image data, a pivot in research focuses towards Pre-CFAR data is observed, aiming to utilize more underlying hidden information.
@@
To our best knowledge, there are currently several datasets containing 4D mmWave radar Pre-CFAR data [65, 64, 41, 77, 78]. Notably, a subset of these datasets [41, 77, 78] have relatively lower elevation resolution, exceeding 15 degrees. Consequently, our survey will focus on the methodologies employed within the high resolution datasets provided by [65, 64]. This section intends to explain the comprehensive pipeline and the enhancements tailored for optimizing 4D mmWave radar characteristics.
@@
As discussed in Section II-A, the 4D mmWave radar signal processing workflow applies FFT methodologies on raw ADC data to discretely process four dimensions: range, Doppler, azimuth, and elevation. This processing generate diverse data representations, including the Range-Doppler (RD) map, Range-Azimuth (RA) map, Range-Azimuth-Doppler (RAD) cube, and ultimately, a 4D tensor. From 2D maps to 4D tensors, the complexity of extracted features varies. Higher dimensional data requires more memory and computation for feature extraction. The extracted features are typically aligned to the RA axis in BEV under polar coordinates or the XY axis in Cartesian coordinates, connecting to detection or segmentation heads, serving as the foundational elements for subsequent detection or segmentation operations.
IV-B1 4D Tensor
Figure 12: Overview of RTNH[72] with 4D mmWave radar tensors
@@
For the 4D tensor, Paek et al. [72, 65] opt to further extract a sparse tensor aligned to the Cartesian coordinate system, subsequently utilizing 3D sparse convolution to extract multi-scale spatial features. Experiments have demonstrated that retaining only the top-5% elements with the highest power measurements can maintain detection accuracy while significantly enhancing processing speed. Furthermore, the elevation information included in 4D tensors is essential facing 3D but not BEV 2D object detection.
IV-B2 RAD Cube
@@ IV-B3 RD Map
IV-B4 RA Map
RA map data, inherently aligned with the polar coordinate system, is amenable to direct processing through dense feature extraction networks. The generated BEV features are then either converted into the Cartesian coordinate system via bi-linear interpolation [41] or utilized within polar-based detection frameworks [71].
@@ IV-B5 Raw ADC Data
IV-C Multi-Modal Fusion Methods
@@
Exemplifying this approach, Meyer et al.[62] adapt a CNN architecture initially designed for camera-LiDAR fusion [79] to process RGB images with height and density maps generated from 4D mmWave radar point clouds. Remarkably, their fusion network demonstrates enhanced precision employing radar instead of LiDAR point clouds, achieving an average precision (AP) of 61% on the Astyx dataset [76]. A subsequent study is performed by Cui et al.[48] with a novel self-supervised model adaptation block[80], which dynamically adapts the fusion of different modalities in accordance with the object properties. Besides, a FV map is generated from the 4D mmWave radar point clouds together with the BEV image. The presented method outperforms the former study[62] by up to 9.5% in 3D AP. The FV map effectively leverages the elevation information provided by 4D mmWave radars and achieves easier fusion with the monocular camera feature, balancing detection accuracy and computational efficiency.
Additionally, recent works such as RCFusion[45] and LXL[58] have advanced the integration of attention mechanisms for the fusion of image and 4D mmWave radar features. They begin by separately extracting BEV features from the camera and radar branches, then employ convolutional networks to create scale-consistent attention maps, effectively delineating the occupancy grid for target objects. The distinction lies in the fact that RCFusion[45] generates 2D attention maps in both the camera and radar branches, which are then multiplied with the BEV feature from the other modality. Conversely, LXL[58] solely utilizes the radar BEV feature to infer the 3D occupancy grid, which is then multiplied with the 3D image voxel features to achieve attention sampling of the image features.
IV-C2 4D Radar with LiDAR
@@
IV-D Challenge
Current methodologies for 4D mmWave radar point cloud perception predominantly adapt established techniques from LiDAR applications, whereas approaches for pre-CFAR data often draw from the vision domain. Although the data formats exhibit similarities, the unique characteristics of mmWave radar data, specifically Doppler velocity and intensity information, warrant more focused attention for effective feature extraction. Moreover, pre-CFAR radar data characteristically contains a significantly higher ratio of background to actual objects (foreground)[69], a factor that complicates data interpretation and model training. The inherent noise within radar data further presents a substantial challenge for learning algorithms. The resilience of 4D mmWave radar models under out-of-distribution (OoD) conditions also remains inadequately explored and understood[65], which underscores the necessity for refined methodologies that can more accurately account for the distinct properties of mmWave radar data, thereby enhancing model robustness and performance in real-world scenarios.
V SLAM Applications
In challenging environments where satellite-based positioning is unreliable or high-definition maps are absent, localization and mapping by perception sensors becomes indispensable. Recently, a collection of SLAM research has been conducted utilizing the emerging 4D mmWave radars.
As Fig. 15 demonstrates, the unique Doppler information contained within radar point clouds presents a notable opportunity for exploitation, which will be discussed in Section V-A. Subsequently, both traditional and learning-based SLAM approaches will be introduced in Section V-B and Section V-C, respectively. And Section V-D briefly introduces current challenges in 4D mmWave radar-based SLAM.
Figure 15: The taxonomy of 4D mmWave radar-based SLAM. Example images are from [7, 81, 4, 82, 83, 84, 85, 86, 87]
V-A Doppler Information Utilizing
The Doppler information constitutes a significant advantage of 4D mmWave radars, especially in the context of SLAM applications. Generally speaking, the utilization of Doppler information can be categorized into several categories:
V-A1 Ego-Velocity Estimation